# Sparse juntas on the biased hypercube

### v4

@@ Abstract
We give a structure theorem for Boolean functions on the $p$-biased hypercube which are $\epsilon$-close to degree $d$ in $L_{2}$, showing that they are close to sparse juntas. Our structure theorem implies that such functions are $O(\epsilon^{C_{d}}+p)$-close to constant functions. We pinpoint the exact value of the constant $C_{d}$.
We also give an analogous result for monotone Boolean functions on the biased hypercube which are $\epsilon$-close to degree $d$ in $L_{2}$, showing that they are close to sparse DNFs.
Our structure theorems are optimal in the following sense: for every $d,\epsilon,p$, we identify a class $\mathcal{F}_{d,\epsilon,p}$ of degree $d$ sparse juntas which are $O(\epsilon)$-close to Boolean (in the monotone case, width $d$ sparse DNFs) such that a Boolean function on the $p$-biased hypercube is $O(\epsilon)$-close to degree $d$ in $L_{2}$ iff it is $O(\epsilon)$-close to a function in $\mathcal{F}_{d,\epsilon,p}$.
@@ 1 Introduction
Let $f\colon\{0,1\}^{n}\to\{0,1\}$ be a Boolean function on the hypercube, where $\{0,1\}^{n}$ is endowed with the uniform measure. If $f$ has degree $1$ then it is a dictator, and if it has degree $d$ then it is a junta [NS94]. Friedgut, Kalai and Naor [FKN02] showed that if $f$ is $\epsilon$-close to degree $1$ (that is, $\|f^{>1}\|^{2}\leq\epsilon$) then $f$ is $O(\epsilon)$-close to a dictator (that is, $\Pr[f\neq g]=O(\epsilon)$ for some dictator $g$). Kindler and Safra [KS02, Kin03] (see also [KK20]) showed that if $f$ is $\epsilon$-close to degree $d$ then $f$ is $O(\epsilon)$-close to a $2^{O(d)}$-junta.
Kindler and Safra proved their result in the more general setting of the biased hypercube, which is $\{0,1\}^{n}$ endowed with the biased measure $\mu_{p}$. They showed that if $f$ is a Boolean function on the $p$-biased hypercube which is $\epsilon$-close to degree $d$ then $f$ is $O(\epsilon+\epsilon^{5/4}/p^{4d})$-close to an $O(1/p^{4d})$-junta, assuming $p\leq 1/2$. When $p$ is constant, this gives the same statement as before, but when $p$ is small, the result is no longer tight.
The biased hypercube in the small $p$ regime is interesting for at least two reasons. First, it is the regime which is most relevant for understanding threshold phenomena in the Erdős–Rényi random graph model $G(n,p)$. Second, it is a good model for other domains with sparse inputs, such as the symmetric group and the Grassmann scheme, the latter of which was recently used to prove the 2-to-2 games conjecture [KMS23].
In this paper, we extend the FKN and Kindler–Safra theorems to functions on the biased hypercube, proving a statement which does not deteriorate as $p$ gets small. For every $d,\epsilon$ and $p\leq 1/2$ we find an “explicit” set of Boolean functions $\mathcal{F}_{d,\epsilon,p}$ such that
@@
    If $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\epsilon$-close to degree $d$ then $f$ is $O(\epsilon)$-close to a function in $\mathcal{F}_{d,\epsilon,p}$.
@@
    Every function in $\mathcal{F}_{d,\epsilon,p}$ is $O(\epsilon)$-close to degree $d$ with respect to $\mu_{p}$.
Let us demonstrate what we mean by explicit using the case $d=1$, which appears in earlier work of Filmus [Fil16]. Filmus showed that we can take
$\mathcal{F}_{1,\epsilon,p}=\left\{\bigvee_{i\in S}y_{i},\lnot\bigvee_{i\in S}y_{i}:|S|\leq\max(1,C\sqrt{\epsilon}/p)\right\},$
@@
The functions in $\mathcal{F}_{1,\epsilon,p}$ are $O(\sqrt{\epsilon})$-close to dictators and so $O(\sqrt{\epsilon}+p)$-close to some constant function, and the same holds for every Boolean function which is $\epsilon$-close to degree $1$. This shows that if we consider Boolean functions which are $\epsilon$-close to degree $1$ up to an error of magnitude $\sqrt{\epsilon}$, then all we see is dictators. If we look closer, narrowing the magnitude of the error to $\epsilon$, then a more nuanced picture emerges. We will be interested in both points of view. We state our results formally in Section 1.1.
When $d>1$, one cannot hope for a set $\mathcal{F}_{d,\epsilon,p}$ which is as explicit as in the case $d=1$. Indeed, even a complete description of all Boolean degree $d$ functions (corresponding to the case $\epsilon=0$) is not available. The set $\mathcal{F}_{d,\epsilon,p}$ that we give is as explicit as we were able to make it; perhaps the reader can improve it. Moreover, it allows us to show that Boolean functions which are $\epsilon$-close to degree $d$ are $O(\epsilon^{C})$-close to Boolean degree $d$ functions, for an appropriate constant $C$, demonstrating the applicability of our main structure theorem.
@@ 1.1 Main results
Here is the generalization of the result of Filmus [Fil16] to arbitrary $d$. The statement uses various terms which we explain after it. The statement is heavily inspired by the FKN theorem for functions on the symmetric group due to Filmus [Fil21].
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\epsilon$-close to degree $d$, where $p\leq 1/2$. Then $f$ is $O(\epsilon)$-close to $\operatorname{round}(g,\{0,1\})$, where $g$ is a degree $d$ polynomial satisfying the following properties, for some constant $C$ depending only on $d$:
@@
    For all $e$, the number of inputs $y\in\{0,1\}^{n}$ of (Hamming) weight $e$ such that $g(y)\notin\{0,1\}$ but $g(z)\in\{0,1\}$ for all $z<y$ is at most $C\epsilon/p^{e}$.
Conversely, if $g$ is a degree $d$ polynomial satisfying these properties then $\operatorname{round}(g,\{0,1\})$ is $O(\epsilon)$-close to degree $d$.
Rather than specifying the list $\mathcal{F}_{d,\epsilon,p}$ directly, we instead specify a list $\mathcal{G}_{d,\epsilon,p}$ of degree $d$ functions. The functions in $\mathcal{F}_{d,\epsilon,p}$ are obtained by rounding the functions in $\mathcal{G}_{d,\epsilon,p}$ to Boolean. The corresponding list for $d=1$ is
$\mathcal{G}_{1,\epsilon,p}=\left\{\sum_{i\in S}y_{i},1-\sum_{i\in S}y_{i}:|S|\leq\max(1,C\sqrt{\epsilon}/p)\right\}.$
A degree $d$ polynomial $g$ is in $\mathcal{G}_{d,\epsilon,p}$ if it satisfies the three properties listed in the theorem. To explain these properties, let us introduce the monomial expansion of $g$, which is its unique expansion as a linear combination of monomials $y_{S}=\prod_{i\in S}y_{i}$. We denote the coefficient of $y_{S}$ by $\tilde{g}(S)$; these are the monomial coefficients of $g$.
The monomial expansion differs from the more usual Fourier expansion, in which instead of the monomials $y_{S}$ we have the Fourier characters $\omega_{S}=\prod_{i\in S}\frac{y_{i}-p}{\sqrt{p(1-p)}}$. Using the monomial expansion guarantees that the functions in $\mathcal{G}_{d,\epsilon,p}$ have the same form regardless of $p$ (cf. $\sum_{i\in S}y_{i}$ to $\sum_{i\in S}\sqrt{p(1-p)}\frac{y_{i}-p}{\sqrt{p(1-p)}}+p|S|$). Concretely, the monomial coefficients appearing in the functions in $\mathcal{G}_{d,\epsilon,p}$ depend only on $d$, a property which fails for the Fourier expansion.
@@
If $\tilde{g}(\emptyset)=0$ then only the first one $y_{1}+y_{2}$ is possible. Defining $m=|\operatorname{supp}(g)|$, the number of minimal non-Boolean inputs is $\binom{m}{2}$, and so Item iii states that $\binom{m}{2}=O(\epsilon/p^{2})$. This implies that either $m\leq 1$ or $m=O(\sqrt{\epsilon}/p)$. The case $\tilde{g}(\emptyset)=1$ is similar.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is a monotone function which is $\epsilon$-close to degree $d$, where $p\leq 1/2$. Then $f$ is $O(\epsilon)$-close to a monotone width $d$ DNF $g$, where $g$ satisfies the following properties, for some constant $C$ depending only on $g$:
@@
    For all $e$, the number of sets $S$ of size $e$ such that $\deg(g|_{S})>d$ but $\deg(g|_{T})\leq d$ for all $T\subsetneq S$ is at most $C\epsilon/p^{e}$.
Conversely, if $g$ is a monotone width $d$ DNF satisfying these properties then it is $O(\epsilon)$-close to degree $d$.
@@
The FKN theorem of Filmus [Fil16] implies that if $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\epsilon$-close to degree $1$ then it is $O(\sqrt{\epsilon})$-close to a dictator and $O(\sqrt{\epsilon}+p)$-close to a constant. Theorem 1.1 implies a similar result for arbitrary $d$, where $\sqrt{\epsilon}$ is replaced by an appropriate power of $\epsilon$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\epsilon$-close to degree $d$, where $p\leq 1/2$. There are parameters $m,M$, depending only on $d$, such that:
@@
    $f$ is $O(\epsilon^{1/M})$-close to a Boolean degree $d$ function.
@@
    $f$ is $O(\epsilon^{1/m}+p)$-close to a Boolean constant function.
@@
$g=y_{1}+\cdots+y_{\sqrt{\epsilon}/p}.$
We can generalize this construction to arbitrary $d$. Let $P$ be a degree $d$ polynomial such that $P(0),\ldots,\allowbreak{P(m-1)}\in\{0,1\}$ and $P(0)\neq P(1)$ (von zur Gathen and Roche [GR97] studied such non-constant polynomials without the restriction $P(0)\neq P(1)$), and take the degree $d$ polynomial
$g=P(y_{1}+\cdots+y_{\epsilon^{1/m}/p}).$
The input to $P$ behaves roughly like a Poisson random variable with expectation $\lambda=\epsilon^{1/m}$, and so $\Pr[g\notin\{0,1\}]=O(\lambda^{m})=O(\epsilon)$. This implies that $\operatorname{round}(g,\{0,1\})$ is $O(\epsilon)$-close to degree $d$.
Since $P(1)\neq P(0)$, a similar calculation shows that $\Pr[g=P(0)]\approx 1-\Theta(\lambda)$ and $\Pr[g=P(1)]\approx\Theta(\lambda)$, and so $g$ is $\Theta(\lambda)$-far from all constant functions.
Using a Ramsey-theoretic argument, we show that this kind of construction is best possible for $m$.
@@
It is an intriguing open question to find the optimal value of $M$. Another interesting exponent is the best $\nu$ such that $f$ is $O(\epsilon^{1/\nu})$-close to some Boolean junta or to some degree $d$ junta (both exponents are the same).
@@ 1.2 Proof sketch
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\epsilon$-close to degree $d$ then there is a Boolean degree $d$ function $g$ which is $O(\epsilon)$-close to $f$.
@@
Suppose we are given a function $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ which is $\epsilon$-close to degree $d$. We look at restrictions $f|_{S}$ obtained by zeroing out the coordinates outside of $S$. If $S\sim\mu_{2p}$ then on average, $f|_{S}$ is close to degree $d$ with respect to the uniform measure on $\{0,1\}^{S}$:
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-f^{\leq d}|_{S})^{2}]\right]=\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\epsilon.$
@@
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-g_{S})^{2}]\right]=O(\epsilon).$ (1)
@@
$\Pr_{\begin{subarray}{c}T\sim\mu_{cq}\\ T_{1},T_{2}\sim\mu_{r}\end{subarray}}[g_{T\cup T_{1}}|_{T}\neq g_{T\cup T_{2}}|_{T}]\leq\epsilon,\text{ where }r=\frac{(1-c)q}{1-cq}.$
@@
The junta agreement theorem follows from the more general agreement theorem in our earlier work [DFH17], which does not require a bound on the size of the monomial support. However, since the proof of the junta agreement theorem is much easier than the proof of the general agreement theorem, we include it in this paper.
@@
$\operatorname*{\mathbb{E}}_{S_{1},S_{2}}[(g_{S_{1}}|_{T}-g_{S_{2}}|_{T})^{2}]=O(\epsilon).$
@@
$\Pr_{S\sim\mu_{p}}[g|_{S}\neq g_{S}]=O(\epsilon).$
A short argument now shows that $\Pr[f\neq g]=O(\epsilon)$.
Since $g$ is obtained by majority decoding and each $g_{S}$ is a junta, we can show that $g$ is a sparse junta, proving Item ii. This allows us to show that $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]=O(\epsilon)$, a step which we highlight below.
@@
In the proof of Theorem 1.1, we deduce the $L_{2}$ guarantee $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]=O(\epsilon)$ from the $L_{0}$ guarantee $\Pr[f\neq g]=O(\epsilon)$ using the fact that $g$ is a sparse junta. The argument uses a technical tool, the reverse union bound, which we would like to highlight here.
Since $\Pr[f\neq g]=O(\epsilon)$ and $f$ is Boolean, it follows that $\Pr[g\notin\{0,1\}]=O(\epsilon)$, and so $\Pr[g(g-1)\neq 0]=O(\epsilon)$. We will deduce that $\operatorname*{\mathbb{E}}_{\mu_{p}}[g^{2}(g-1)^{2}]=O(\epsilon)$, which implies $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]=O(\epsilon)$ via a short argument.
Since $g$ is a sparse junta whose monomial coefficients are quantized, the same holds for $h=g(g-1)$, and also for $h^{2}$. We will show that $\Pr[h\neq 0]=O(\epsilon)$ implies that $\operatorname*{\mathbb{E}}_{\mu_{p}}[h^{2}]=O(\epsilon)$.
Since the monomial coefficients of $h^{2}$ are quantized, it suffices to show that the monomial support of $h^{2}$ is sparse, in the sense that it contains $O(\epsilon/p^{e})$ sets of size $e$ for all $e\leq 2d$. Since $h$ is a sparse junta, in order to show that the monomial support of $h^{2}$ is sparse, it suffices to show that the monomial support of $h$ itself is sparse.
@@
If the inequality were in the other direction then it would follow that the monomial support of $h$ is sparse, since $\Pr[h\neq 0]=O(\epsilon)$. In general, we cannot reverse the inequality (consider for example $h=\sum_{i=1}^{n}y_{i}$, where the left-hand side tends to $1$ while the right-hand side tends to infinity for constant $p$). However, when $h$ is a sparse junta, the reverse union bound states that we can indeed reverse the direction of the inequality:
@@
This paper originates in an extended abstract [DFH19], which covers both the material eventually published in [DFH17] and part of the material in the present work. The full version of [DFH19] is available on arXiv (1711.09428) and ECCC (TR17-180).
The full version of [DFH19] contains weaker versions of Theorems 1.1 and 1.2, which describe only some properties of the approximating function $g$, and in particular do not constitute a full characterization of all Boolean functions which are close to degree $d$. The current versions of these theorems are inspired by [Fil21]. In addition, Theorem 1.4 is completely new. It is inspired by [Fil23].
In contrast, the main theorems in the full version of [DFH19] apply in the more general $A$-valued setting, which we describe briefly in Section 7, as well as in the setting of the slice. We omit these generalizations in this version to make it more concise, and since these generalizations only require a few new ideas, the most interesting of which is that whereas the monomial expansion of functions on the slice is not unique, the sparse monomial expansion of a junta is unique.
@@
Acknowledgements
We thank the anonymous reviewers for their careful reading of the manuscript, and for significantly improving some of the proofs.

### v6

@@ Abstract
We give a structure theorem for Boolean functions on the $p$-biased hypercube which are $\varepsilon$-close to degree $d$ in $L_{2}$, showing that they are close to sparse juntas. Our structure theorem implies that such functions are $O(\varepsilon^{C_{d}}+p)$-close to constant functions. We pinpoint the exact value of the constant $C_{d}$.
We also give an analogous result for monotone Boolean functions on the biased hypercube which are $\varepsilon$-close to degree $d$ in $L_{2}$, showing that they are close to sparse DNFs.
Our structure theorems are optimal in the following sense: for every $d,\varepsilon,p$, we identify a class $\mathcal{F}_{d,\varepsilon,p}$ of degree $d$ sparse juntas which are $O(\varepsilon)$-close to Boolean (in the monotone case, width $d$ sparse DNFs) such that a Boolean function on the $p$-biased hypercube is $O(\varepsilon)$-close to degree $d$ in $L_{2}$ if and only if it is $O(\varepsilon)$-close to a function in $\mathcal{F}_{d,\varepsilon,p}$.
@@ 1 Introduction
Let $f\colon\{0,1\}^{n}\to\{0,1\}$ be a Boolean function on the hypercube, where $\{0,1\}^{n}$ is endowed with the uniform measure. If $f$ has degree $1$ then it is a dictator, and if it has degree $d$ then it is a junta [NisanS1994]. Friedgut, Kalai, and Naor (FKN)  [FriedgutKN2002] showed that if $f$ is $\varepsilon$-close to degree $1$ (that is, $\|f^{>1}\|^{2}\leq\varepsilon$) then $f$ is $O(\varepsilon)$-close to a dictator (that is, $\Pr[f\neq g]=O(\varepsilon)$ for some dictator $g$). Kindler and Safra [KindlerS2002, Kindler2003] (see also [KellerK2020]) showed that if $f$ is $\varepsilon$-close to degree $d$ then $f$ is $O(\varepsilon)$-close to a $2^{O(d)}$-junta.
Kindler and Safra proved their result in the more general setting of the biased hypercube, which is $\{0,1\}^{n}$ endowed with the biased measure $\mu_{p}$. They showed that if $f$ is a Boolean function on the $p$-biased hypercube which is $\varepsilon$-close to degree $d$ then $f$ is $O(\varepsilon+\varepsilon^{5/4}/p^{4d})$-close to an $O(1/p^{4d})$-junta, assuming $p\leq 1/2$. When $p$ is constant, this gives the same statement as before, but when $p$ is small, the result is no longer tight.
The biased hypercube in the small $p$ regime is interesting for at least two reasons. First, it is the regime that is most relevant for understanding threshold phenomena in the Erdős–Rényi random graph model $G(n,p)$. Second, it is a good model for other domains with sparse inputs, such as the symmetric group and the Grassmann scheme, the latter of which was recently used to prove the 2-to-2 games conjecture [KhotMS2023].
In this paper, we extend the FKN and Kindler–Safra theorems to functions on the biased hypercube, proving a statement which does not deteriorate as $p$ gets small. For every $d,\varepsilon$ and $p\leq 1/2$ we find an “explicit” set of Boolean functions $\mathcal{F}_{d,\varepsilon,p}$ (consisting of degree $d$ “sparse juntas”, a concept we define below) such that
@@
    If $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\varepsilon$-close to degree $d$ then $f$ is $O(\varepsilon)$-close to a function in $\mathcal{F}_{d,\varepsilon,p}$.
@@
    Every function in $\mathcal{F}_{d,\varepsilon,p}$ is $O(\varepsilon)$-close to degree $d$ with respect to $\mu_{p}$.
Let us demonstrate what we mean by explicit using the case $d=1$, which appears in earlier work of Filmus [Filmus2016-fkn]. Filmus showed that we can take
$\mathcal{F}_{1,\varepsilon,p}=\left\{\bigvee_{i\in S}y_{i},\lnot\bigvee_{i\in S}y_{i}:|S|\leq\max(1,C\sqrt{\varepsilon}/p)\right\},$
@@
The functions in $\mathcal{F}_{1,\varepsilon,p}$ are $O(\sqrt{\varepsilon})$-close to dictators and so $O(\sqrt{\varepsilon}+p)$-close to some constant function, and the same holds for every Boolean function which is $\varepsilon$-close to degree $1$. This shows that if we consider Boolean functions which are $\varepsilon$-close to degree $1$ up to an error of magnitude $\sqrt{\varepsilon}$, then all we see is dictators. If we look closer, narrowing the magnitude of the error to $\varepsilon$, then a more nuanced picture emerges. We will be interested in both points of view. We state our results formally in Section 1.1.
When $d>1$, one cannot hope for a set $\mathcal{F}_{d,\varepsilon,p}$ which is as explicit as in the case $d=1$. Indeed, even a complete description of all Boolean degree $d$ functions (corresponding to the case $\varepsilon=0$) is not available. The set $\mathcal{F}_{d,\varepsilon,p}$ that we give is as explicit as we were able to make it; perhaps the reader can improve it. Moreover, it allows us to show that Boolean functions which are $\varepsilon$-close to degree $d$ are $O(\varepsilon^{C})$-close to Boolean degree $d$ functions, for an appropriate constant $C$, demonstrating the applicability of our main structure theorem.
@@ 1.1 Main results
Here is the generalization of the result of Filmus [Filmus2016-fkn] to arbitrary $d$. The statement uses various terms which we explain after it. The statement is heavily inspired by the FKN theorem for functions on the symmetric group due to Filmus [Filmus2021].
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Then $f$ is $O(\varepsilon)$-close to $\operatorname{round}(g,\{0,1\})$, where $g$ is a degree $d$ polynomial satisfying the following properties, for some constant $C$ depending only on $d$:
@@
    For all $e$, the number of inputs $y\in\{0,1\}^{n}$ of (Hamming) weight $e$ such that $g(y)\notin\{0,1\}$ but $g(z)\in\{0,1\}$ for all $z<y$ is at most $C\varepsilon/p^{e}$.
Conversely, if $g$ is a degree $d$ polynomial satisfying these properties then $\operatorname{round}(g,\{0,1\})$ is $O(\varepsilon)$-close to degree $d$.
Rather than specifying the list $\mathcal{F}_{d,\varepsilon,p}$ directly, we instead specify a list $\mathcal{G}_{d,\varepsilon,p}$ of degree $d$ functions. The functions in $\mathcal{F}_{d,\varepsilon,p}$ are obtained by rounding the functions in $\mathcal{G}_{d,\varepsilon,p}$ to Boolean. The corresponding list for $d=1$ is
$\mathcal{G}_{1,\varepsilon,p}=\left\{\sum_{i\in S}y_{i},1-\sum_{i\in S}y_{i}:|S|\leq\max(1,C\sqrt{\varepsilon}/p)\right\}.$
A degree $d$ polynomial $g$ is in $\mathcal{G}_{d,\varepsilon,p}$ if it satisfies the three properties listed in the theorem. To explain these properties, let us introduce the monomial expansion of $g$, which is its unique expansion as a linear combination of monomials $y_{S}=\prod_{i\in S}y_{i}$. We denote the coefficient of $y_{S}$ by $\tilde{g}(S)$; these are the monomial coefficients of $g$.
The monomial expansion differs from the more usual Fourier expansion, in which instead of the monomials $y_{S}$ we have the Fourier characters $\omega_{S}=\prod_{i\in S}\frac{y_{i}-p}{\sqrt{p(1-p)}}$. Using the monomial expansion guarantees that the functions in $\mathcal{G}_{d,\varepsilon,p}$ have the same form regardless of $p$ (cf. $\sum_{i\in S}y_{i}$ to $\sum_{i\in S}\sqrt{p(1-p)}\frac{y_{i}-p}{\sqrt{p(1-p)}}+p|S|$). Concretely, the monomial coefficients appearing in the functions in $\mathcal{G}_{d,\varepsilon,p}$ depend only on $d$, a property which fails for the Fourier expansion.
@@
If $\tilde{g}(\emptyset)=0$ then only the first one $y_{1}+y_{2}$ is possible. Defining $m=|\operatorname{supp}(g)|$, the number of minimal non-Boolean inputs is $\binom{m}{2}$, and so Item iii states that $\binom{m}{2}=O(\varepsilon/p^{2})$. This implies that either $m\leq 1$ or $m=O(\sqrt{\varepsilon}/p)$. The case $\tilde{g}(\emptyset)=1$ is similar.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is a monotone function which is $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Then $f$ is $O(\varepsilon)$-close to a monotone width $d$ DNF $g$, where $g$ satisfies the following properties, for some constant $C$ depending only on $g$:
@@
    For all $e$, the number of sets $S$ of size $e$ such that $\deg(g|_{S})>d$ but $\deg(g|_{T})\leq d$ for all $T\subsetneq S$ is at most $C\varepsilon/p^{e}$.
Conversely, if $g$ is a monotone width $d$ DNF satisfying these properties then it is $O(\varepsilon)$-close to degree $d$.
@@
The FKN theorem of Filmus [Filmus2016-fkn] implies that if $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\varepsilon$-close to degree $1$ then it is $O(\sqrt{\varepsilon})$-close to a dictator and $O(\sqrt{\varepsilon}+p)$-close to a constant. Theorem 1.1 implies a similar result for arbitrary $d$, where $\sqrt{\varepsilon}$ is replaced by an appropriate power of $\varepsilon$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\varepsilon$-close to degree $d$, where $p\leq 1/2$. There are parameters $m,M$, depending only on $d$, such that:
@@
    $f$ is $O(\varepsilon^{1/M})$-close to a Boolean degree $d$ function.
@@
    $f$ is $O(\varepsilon^{1/m}+p)$-close to a Boolean constant function.
@@
$g=y_{1}+\cdots+y_{\sqrt{\varepsilon}/p}.$
We can generalize this construction to arbitrary $d$. Let $P$ be a degree $d$ polynomial such that $P(0),\ldots,\allowbreak{P(m-1)}\in\{0,1\}$ and $P(0)\neq P(1)$ (von zur Gathen and Roche [GathenR1997] studied such non-constant polynomials without the restriction $P(0)\neq P(1)$), and take the degree $d$ polynomial
$g=P(y_{1}+\cdots+y_{\varepsilon^{1/m}/p}).$
The input to $P$ behaves roughly like a Poisson random variable with expectation $\lambda=\varepsilon^{1/m}$, and so $\Pr[g\notin\{0,1\}]=O(\lambda^{m})=O(\varepsilon)$. This implies that $\operatorname{round}(g,\{0,1\})$ is $O(\varepsilon)$-close to degree $d$.
Since $P(1)\neq P(0)$, a similar calculation shows that $\Pr[g=P(0)]\approx 1-\Theta(\lambda)$ and that $\Pr[g=P(1)]\approx\Theta(\lambda)$, and so $g$ is $\Theta(\lambda)$-far from all constant functions.
Using a Ramsey-theoretic argument, we show that this kind of construction is the best possible for $m$.
@@
It is an intriguing open question to find the optimal value of $M$. Another interesting exponent is the best $\nu$ such that $f$ is $O(\varepsilon^{1/\nu})$-close to some Boolean junta or to some degree $d$ junta (both exponents are the same).
@@ 1.2 Proof sketch
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\varepsilon$-close to degree $d$ then there is a Boolean degree $d$ function $g$ which is $O(\varepsilon)$-close to $f$.
@@
Suppose we are given a function $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ which is $\varepsilon$-close to degree $d$. We look at restrictions $f|_{S}$ obtained by zeroing out the coordinates outside of $S$. If $S\sim\mu_{2p}$ then on average, $f|_{S}$ is close to degree $d$ with respect to the uniform measure on $\{0,1\}^{S}$:
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-f^{\leq d}|_{S})^{2}]\right]=\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\varepsilon.$
@@
$\operatorname*{\mathbb{E}}_{S\sim\mu_{2p}}\left[\operatorname*{\mathbb{E}}_{\mu_{1/2}}[(f|_{S}-g_{S})^{2}]\right]=O(\varepsilon).$ (1)
@@
$\Pr_{\begin{subarray}{c}T\sim\mu_{cq}\\ T_{1},T_{2}\sim\mu_{r}\end{subarray}}[g_{T\cup T_{1}}|_{T}\neq g_{T\cup T_{2}}|_{T}]\leq\varepsilon,\text{ where }r=\frac{(1-c)q}{1-cq}.$
@@
The junta agreement theorem follows from the more general agreement theorem in our earlier work [DinurFH-agree], which does not require a bound on the size of the monomial support. However, since the proof of the junta agreement theorem is much easier than the proof of the general agreement theorem, we include it in this paper.
@@
$\operatorname*{\mathbb{E}}_{S_{1},S_{2}}[(g_{S_{1}}|_{T}-g_{S_{2}}|_{T})^{2}]=O(\varepsilon).$
@@
$\Pr_{S\sim\mu_{p}}[g|_{S}\neq g_{S}]=O(\varepsilon).$
A short argument now shows that $\Pr[f\neq g]=O(\varepsilon)$.
Since $g$ is obtained by majority decoding and each $g_{S}$ is a junta, we can show that $g$ is a sparse junta, proving Item ii. This allows us to show that $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]=O(\varepsilon)$, a step which we highlight below.
@@
In the proof of Theorem 1.1, we deduce the $L_{2}$ guarantee $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]=O(\varepsilon)$ from the $L_{0}$ guarantee $\Pr[f\neq g]=O(\varepsilon)$ using the fact that $g$ is a sparse junta. The argument uses a technical tool, the reverse union bound, which we would like to highlight here.
Since $\Pr[f\neq g]=O(\varepsilon)$ and $f$ is Boolean, it follows that $\Pr[g\notin\{0,1\}]=O(\varepsilon)$, and so $\Pr[g(g-1)\neq 0]=O(\varepsilon)$. We will deduce that $\operatorname*{\mathbb{E}}_{\mu_{p}}[g^{2}(g-1)^{2}]=O(\varepsilon)$, which implies $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]=O(\varepsilon)$ via a short argument.
Since $g$ is a sparse junta whose monomial coefficients are quantized, the same holds for $h=g(g-1)$, and also for $h^{2}$. We will show that $\Pr[h\neq 0]=O(\varepsilon)$ implies that $\operatorname*{\mathbb{E}}_{\mu_{p}}[h^{2}]=O(\varepsilon)$.
Since the monomial coefficients of $h^{2}$ are quantized, it suffices to show that the monomial support of $h^{2}$ is sparse, in the sense that it contains $O(\varepsilon/p^{e})$ sets of size $e$ for all $e\leq 2d$. Since $h$ is a sparse junta, in order to show that the monomial support of $h^{2}$ is sparse, it suffices to show that the monomial support of $h$ itself is sparse.
@@
If the inequality were in the other direction then it would follow that the monomial support of $h$ is sparse, since $\Pr[h\neq 0]=O(\varepsilon)$. In general, we cannot reverse the inequality (consider for example $h=\sum_{i=1}^{n}y_{i}$, where the left-hand side tends to $1$ while the right-hand side tends to infinity for constant $p$). However, when $h$ is a sparse junta, the reverse union bound states that we can indeed reverse the direction of the inequality:
@@
This paper originates in an extended abstract [DinurFH2019], which covers both the material eventually published in [DinurFH-agree] and part of the material in the present work. The full version of [DinurFH2019] is available on arXiv (1711.09428) and ECCC (TR17-180).
The full version of [DinurFH2019] contains weaker versions of Theorems 1.1 and 1.2, which describe only some properties of the approximating function $g$, and in particular do not constitute a full characterization of all Boolean functions which are close to degree $d$. The current versions of these theorems are inspired by [Filmus2021]. In addition, Theorem 1.4 is completely new. It is inspired by [Filmus2023].
In contrast, the main theorems in the full version of [DinurFH2019] apply in the more general $A$-valued setting, which we describe briefly in Section 7, as well as in the setting of the slice. We omit these generalizations in this version to make it more concise, and since these generalizations only require a few new ideas, the most interesting of which is that whereas the monomial expansion of functions on the slice is not unique, the sparse monomial expansion of a junta is unique.